Sove each system by substitution method. I do not understand this at all, can someone help with explanations, Thanks.
x+3y=2
-x+y=1

Answer by checkley77(12844)

x+3y=2
x=2-3y
Now replace x in the second equation with (2-3y)
-x+y=1
-(2-3y)+y=1
-2+3y+y=1
4y=1+2
4y=3
y=3/4 ans.
x=2-3*3/4
x=2-9/4
x=(2*4-9)/4
x=(8-9)/4
x=-1/4 ans.
Proof:
1(-1/4)+3/4==1
1/4+3/4=1
1=1